OpenStudy (anonymous):

What is the solution to the equation log(base 4)x – log(base 4) 2 = 2 What is the solution to the equation log(base 4)x – log(base 4) 2 = 2 @Mathematics

OpenStudy (alfie):

\[\log_4x - \log_42 = 2 \rightarrow \log_4\frac{x}{2} = 2 \rightarrow 4^2 = \frac{x}{2} \rightarrow x = 32\]